A golf club with 70j of energy strikes a stationary golf ball with a mass of 48.0 g. The energy transfer is only 21% efficientCalculate the initial speed of the golf ball. 


1
Expert's answer
2021-01-12T12:13:37-0500

Let's first find the kinetic energy transfered to golf ball:


"KE_{ball}=0.21KE_{golf\ club}=0.21\cdot70\ J=14.7\ J."

Finally, we can find the initial speed of the golf ball after the hitting by the golf club:


"KE_{ball}=\dfrac{1}{2}mv^2,""v=\sqrt{\dfrac{2KE_{ball}}{m}}=\sqrt{\dfrac{2\cdot 14.7\ J}{4.8\cdot10^{-2}\ kg}}=24.75\ \dfrac{m}{s}."

Answer:

"v=24.75\ \dfrac{m}{s}."